Grade 11General Physics

find the moment of inertia of a rod of mass m and length l about an axis which passes from one end of the rod and makes angle theta with the length of the rod

riya
10 Years agoGrade 11

1 Answer

Riddhish Bhalodia
10 Years ago
The easiest way to do this to break the rod into components along and perpendicular to the axis

The along component will have the lenght lcos(theta) but will have moment of inertia =0
The moment of inertia is only by the perpendicular component usually given by
I = 1/3ML^2
here
L = lsin\theta, \quad M = msin\theta
so
I = 1/3ml^2sin^3\theta